### v2

@@ Abstract
Social media platforms systematically reward popularity over authenticity, incentivizing users to strategically tailor their expression for attention. In this paper, we introduce (i) an attention-seeking model, distinct from canonical mechanisms of conformity, learning, persuasion, and (mis)information transmission in social networks literature, and (ii) the first utilitarian framework defined directly over observable social media platform metrics, filling a critical gap in the social media literature. In the model, agents hold fixed heterogeneous authentic opinions and derive (i) utility gains from the popularity of their own posts—measured by likes received, and (ii) utility gains (losses) from exposure to content that aligns with (diverges from) their authentic opinion. Social media interaction acts as a state-dependent welfare amplifier: light topics generate Pareto improvements, whereas intense topics make everyone worse off in a polarized society (e.g., political debates during elections). Moreover, strategic expression amplifies social media polarization during polarized events while dampening it during unified events (e.g., national celebrations). Consequently, strategic distortions magnify welfare outcomes, expanding aggregate gains in light topics while exacerbating losses in intense, polarized ones. Counterintuitively, strategic agents often face a popularity trap: posting a more popular opinion is individually optimal, yet collective action by similar agents eliminates their authentic opinion from the platform, leaving them worse off than under the authentic-expression benchmark. Preference-based algorithms—widely used by platforms— or homophilic exposures discipline popularity-driven behavior, narrowing the popularity trap region and limiting its welfare effects.
@@ 1  Introduction
Social media platforms are defined by a distinctive feature: the quantification of social approval through “likes”. This mechanism is often tied to monetary or non-monetary rewards ([23], [18], [42], [6], [24]), thereby incentivizing users to seek popularity for their content. When applied to opinion expression, this feature alters the strategic landscape: incentivizing agents to trade off popularity against authenticity.
In this paper, we study strategic expression in social media. We show that social media acts as a state-dependent welfare amplifier. Moreover, popularity-driven strategic expression systematically reshapes the landscape of online discourse, generating redistributive and/or magnifying effects on polarization, individual utilities, and aggregate welfare.
@@
Our main contribution is twofold. First, our model introduces popularity (or attention-seeking) as an unexplored channel in the social networks literature, distinct from the canonical mechanisms of conformity, learning, belief updating, and persuasion (e.g., [7, 21, 10, 36, 25, 28, 33, 35, 11, 29]) [^2] Readers are referred to [39] for a model of strategic information sharing in social media with learning in a voting framework.. In our analysis, we abstract away from these canonical mechanisms to discover this new strategic landscape, shedding light on a potential disconnect between beliefs and online discourse. Relatedly, empirical studies show that social media often has limited effects on beliefs ([27, 13, 38]).
@@
Second, we are the first to introduce a utilitarian framework defined directly over social media metrics: posts and likes. As emphasized in the review by [6], the existing literature evaluates welfare indirectly—through access choices, time use, or subjective well-being (e.g., [3, 4, 15, 14])—rather than specifying utility directly over observable metrics. Our tractable utilitarian framework fills this critical gap in the social media literature. This workhorse model allows us to measure individual utilities and aggregate welfare directly from social media interaction.333[6] emphasize in their review: …While most existing work has concentrated on the time spent on social media, the content consumed and the nature of the engagement are also important. Existing work has primarily quantified welfare through stated-preference measures, such as willingness to pay or accept for data sharing and privacy, which face well-known challenges including uncertainty, externalities, and gaps between stated and revealed preferences. Our framework directly addresses this gap while isolating from learning/persuasion channels. Developing a hybrid microfoundation that combines popularity-driven incentives with learning/persuasion is a natural direction for future research.
@@
In our model, agents hold heterogeneous and fixed authentic opinions, yet derive utility from social media activities: (i) utility gains from the popularity of their own posts, and (ii) utility gains (losses) from exposure to content that aligns with (diverges from) their authentic opinion. To purely isolate the popularity-driven expression, we first focus on a representative social media environment in which each agent’s audience and exposure mirror the societal distribution of opinions. This benchmark abstracts from friendship and/or algorithmic biases—such as homophily—and clearly identifies the mechanism. In Section 6, we extend our analysis to algorithmic filtering by modeling a targeting rule widely used across social media platforms.
@@
Societal opinions are often sharply polarized, from intense political debates to light, taste-based topics, or sharply unified around broad consensus (e.g., national celebrations or responses to natural disasters). We apply our framework to these environments by introducing a tractable three-opinion space. In particular, we define an opinion space where each agent holds an authentic opinion in the set $\{-b,0,b\}$, a polarized or a neutral (central) view. The magnitude $|b|$ measures the topic intensity—representing the intensity of opposition between opinions—which serves as a key driver of strategic actions and resulting utility gains and losses.
@@
We show a fundamental duality in how social media polarization differs from societal polarization: When society is already divided (e.g., political debates, immigration policy debates), popularity incentives amplify polarization in social media by inducing neutral agents to post opinionated content. Conversely, when society is highly unified (e.g., national celebrations, national sports events), the same incentives drive opinionated agents to converge toward neutral expression, generating amplified online unification. Social media thus acts as a non-linear magnifier, selectively distorting the perception of societal division depending on the nature of the event. Unlike existing work that studies online polarization (e.g., [27, 8, 9, 20, 39, 19, 31, 22, 41, 2, 12]), our results focus on the role of popularity-driven strategic expression in it. Among these studies, [20] analyze polarization in a social media network with two news content types—mass-market and niche—where users differ in their preferred content to recommend. In their model, individuals recommend their preferred type of content when it is available in their network, and otherwise recommend the alternative one. They show that when the friendship network is not biased, content filtering tends to amplify mass-market content and reduce polarization at steady state— defined as the difference between the mass and niche content recommendation probabilities. However, greater network connectivity and homophily may increase the prevalence of niche content and, in turn, increase polarization. In another related work, [39] analyzes content filtering in a similar two-type model within a voting framework. In their model, agents receive private signals and strategically share content based on their own partisan views and the composition of their audience on social media. Sharing decisions are strategic, whereas voting decisions are sincere and reflect Bayesian updating of beliefs about the state of the world. Different from these two models, our framework, in a multiple opinion setup, allows agents to strategically choose what to post (from a continuous opinion space) based on popularity and alignment incentives. Moreover, in a three-opinion setting, we show that the impact of strategic expression on polarization is state-dependent: polarization is amplified (dampened) when the underlying society is highly polarized (less polarized). Unlike [20], network density plays no role in shaping polarization in our model, though it has sharp implications for user welfare, a dimension not addressed in these two studies.
@@
In this framework with sticky opinions and popularity incentives, the welfare consequences are sharply state-dependent. For light topics, social media is welfare-improving because the utility gains from popularity and exposure to aligned content exceeds the utility losses from exposure to misaligned yet low-stakes content. However, for intense topics, social media interaction is often harmful for some groups of agents; and for high-polarization events, intense topics create an environment in which everyone is worse off from social media interaction due to high misalignment costs. We also show that network density plays a purely welfare-amplifying role: denser networks increase exposure intensity, without altering equilibrium posting behavior. As a result, higher density magnifies welfare gains when social media interaction is welfare-improving, and amplifies welfare losses when it is welfare-reducing.
@@
Crucially, in both high- and low-polarization cases, strategic posting behavior has heterogeneous effects for different opinion holders. Relatedly, we identify a phenomenon that we call the “Popularity Trap”: for intense topics, social media activity can be utility-reducing for agents who strategically post a more popular opinion. Individually-optimal deviations by each such member cause their authentic opinion to disappear from social media, which amplifies their exposure to misaligned content, creating an equilibrium outcome where these strategic agents are worse off compared to the authentic expression benchmark. The popularity trap applies for neutral (opinionated) agents during high-(low-)polarization events.
@@
Relatedly, our framework is distinct from the widely studied conformity models (e.g., [7, 10, 36, 43, 16]). In conformity models, utility typically increases with similarity to others’ actions. These incentives generate coordination benefits: an agent benefits when others also conform to the same norm. In our framework, the direction of externalities is reversed. Agents do not desire collective convergence; rather, they prefer to enjoy higher utility from popularity, but prefer that other agents who share their authentic opinion continue to express it authentically. This generates distinct utility and welfare implications compared to conformity models, and also introduces the popularity traps.
Finally, our framework departs from existing models of content transmission online—whether through network diffusion (e.g., [5, 41, 40], [1]), algorithmic filtering (e.g., [34, 30, 8]), selective sharing (e.g., [20, 39, 26]), or spread of misinformation (e.g., [37, 32, 2])—and instead models it as a popularity-driven environment.
@@ 3  Expressed Opinions and Distorted Utilities
Lemma 1999The formal proof of Lemma 1 and the proof of Proposition 6 are in the Online Appendix. relies on a duality in equilibrium: agents’ authentic preferences are reflected in their likes on others’ posts but not necessarily in their own posts. At equilibrium, agent $i$ likes agent $j$’s post if and only if the post aligns with her own opinion, $c_{j}=b_{i}$. For any $c_{j}\neq b_{i}$, liking $j$’s post yields disutility $\omega^{d}_{i}|b_{i}-c_{j}|$, and therefore $i$ chooses not to react ($r_{ij}=0$). In contrast, when $j$’s post aligns with $i$’s opinion ($c_{j}=b_{i}$), liking provides a positive payoff $\omega^{a}_{i}$. Hence, $r_{ij}=1$ if and only if $c_{j}=b_{i}$.
@@ 4  Polarization and Welfare on Social Media
The equilibrium group sizes representing opinions b-b, 0, and +b+b on social media are denoted by $C_{-}$, $C_{0}$, and $C_{+}$, respectively. Following Proposition 1, $C_{-}+C_{0}+C_{+}=n$ holds. To avoid divisibility issues, we assume that $n$ and $G_{0}$ are even numbers. Moreover, if there exists some neutral agents who are indifferent between posting one of the opinionated views b-b and +b+b, we assume that the number of such indifferent neutral agents is always even, and the equilibrium posts of these agents are evenly split between $C_{-}$ and $C_{+}$.

### v3

@@ Abstract
Social media platforms systematically reward popularity over authenticity, incentivizing users to strategically tailor their expression for attention. In this paper, we introduce (i) popularity as a strategic expression mechanism, distinct from the canonical mechanisms of conformity, learning, persuasion, and (mis)information transmission in social networks, and (ii) a utilitarian framework for measuring user welfare that maps directly to observable platform metrics, filling a critical gap in the social media literature. In the model, agents hold fixed heterogeneous authentic opinions and derive utility gains from the popularity of their own posts—measured by likes received, and utility gains (losses) from exposure to content that aligns with (diverges from) their authentic opinion. Social media interaction acts as a state-dependent welfare amplifier: light topics generate Pareto improvements, whereas intense topics make everyone worse off in a polarized society (e.g., political debates during elections). Moreover, strategic expression amplifies social media polarization during polarized events while dampening it during unified events (e.g., national celebrations). Consequently, strategic distortions magnify welfare outcomes, expanding aggregate gains in light topics while exacerbating losses in intense, polarized ones. Counterintuitively, strategic agents often face a popularity trap: posting a more popular opinion is individually optimal, yet collective action by similar agents eliminates their authentic opinion from the platform, leaving them worse off than under the authentic-expression benchmark. Homophilic algorithms that match users with preferred content—widely used by platforms—discipline popularity-driven behavior, narrowing the popularity trap region and limiting its welfare effects.
@@ 1  Introduction
Social media platforms are defined by a distinctive feature: the quantification of social approval through “likes.” This mechanism is often tied to monetary or non-monetary rewards ([25], [19], [45], [6], [27]), thereby incentivizing users to seek popularity—measured by likes received—of their content. When applied to opinion expression, this feature alters the strategic landscape by incentivizing agents to trade off popularity against authenticity. How does this strategic tension shape online polarization and user welfare from social media interaction?
We develop a model of strategic expression in social media and show that it generates redistributive and/or magnifying effects on social media polarization, individual utilities, and aggregate welfare. Moreover, homophilic algorithms widely used by platforms discipline rather than amplify these popularity-driven distortions, suggesting that targeted content delivery acts as a moderating force on strategic expression—and, contrary to common belief, may temper rather than fuel online polarization.
@@
Our main contribution is twofold. First, we introduce an unexplored strategic expression channel in social networks, distinct from the canonical mechanisms of conformity, learning, belief updating, and persuasion (e.g., [7, 23, 10, 39, 28, 30, 36, 38, 12, 42, 26, 22]) [^2] Readers are referred to [42] for a model of strategic information sharing in social media with learning in a voting framework.. In our model, agents hold heterogeneous and fixed authentic opinions, and derive utility gains from the popularity of their own posts, and utility gains (losses) from exposure to content that aligns with (diverges from) their authentic opinion, featuring a trade-off between authenticity and popularity as well as externalities from others’ content. This framework allows us to abstract away from the canonical mechanisms above and discover a new strategic landscape, shedding light on a disconnect between beliefs and online discourse. Relatedly, empirical studies by [29], [14], and [41] show that social media often has limited effects on beliefs. Besides, while a rich literature in behavioral economics documents how revealed preferences diverge from normative preferences—with [11] identifying the key factors driving this divergence—our model introduces a new such factor that specifically applies to online interaction in social networks: the pursuit of popularity. In this sense, our framework departs from existing models of content transmission online—whether through network diffusion (e.g., [5, 44, 43], [1]), algorithmic filtering (e.g., [37, 33, 8]), or spread of misinformation (e.g., [40, 35, 2])—and instead models it as a strategic expression environment.
@@
Second, we are the first to introduce a utilitarian framework for welfare analysis in social media that is defined directly over social media metrics: posts and likes. As emphasized in the review by [6], the existing literature evaluates welfare indirectly—through access choices, time use, or subjective well-being (e.g., [3, 4, 16, 15])—rather than specifying utility directly over observable metrics. Our tractable utilitarian framework fills this critical gap in the social media literature, allowing us to measure individual utilities and aggregate welfare directly from social media interaction.333[6] emphasize in their review: …While most existing work has concentrated on the time spent on social media, the content consumed, and the nature of the engagement are also important. Existing work has primarily quantified welfare through stated-preference measures, such as willingness to pay or accept for data sharing and privacy, which face well-known challenges including uncertainty, externalities, and gaps between stated and revealed preferences. Our framework directly addresses this gap while isolating from learning/persuasion channels. Developing a hybrid microfoundation that combines popularity-driven incentives with learning/persuasion is a natural direction for future research.
@@
In our analysis, we first focus on a representative social media environment in which each agent’s audience and exposure mirror the societal distribution of opinions. This benchmark abstracts from friendship and/or algorithmic biases—such as homophily—and clearly identifies the underlying mechanism. In Section 6, we extend our analysis to algorithmic filtering by modeling a targeting rule widely used across social media platforms.
@@
Societal opinions are often sharply polarized, from intense political debates to light, taste-based topics, or sharply unified around broad consensus (e.g., national celebrations or responses to natural disasters). We apply our framework to these environments by introducing a tractable three-opinion space. In particular, we define an opinion space where each agent holds an authentic opinion in the set $\{-b,0,b\}$, a polarized or neutral (central) view. The magnitude $|b|$ measures the topic intensity—representing the intensity of opposition between opinions—which serves as a key driver of strategic actions and resulting utility gains and losses.
@@
We show a fundamental duality in how social media polarization differs from societal polarization: When society is already divided (e.g., political debates, immigration policy debates), popularity incentives amplify polarization in social media by inducing neutral agents to post opinionated content. Conversely, when society is highly unified (e.g., national celebrations, national sports events), the same incentives drive opinionated agents to converge toward neutral expression, generating amplified online unification. Social media thus acts as a non-linear magnifier, selectively distorting the perception of societal division depending on the nature of the event. Unlike existing work that studies online polarization (e.g., [29, 8, 9, 21, 42, 20, 34, 24, 44, 2, 13]), our results focus on the role of popularity-driven strategic expression in it. Among these studies, [21] analyze polarization in a social media network with two news content types—mass-market and niche—where users differ in their preferred content to recommend. In their model, individuals recommend their preferred type of content when it is available in their network, and otherwise recommend the alternative one. They show that when the friendship network is not biased, content filtering tends to amplify mass-market content and reduce polarization at steady state— defined as the difference between the mass and niche content recommendation probabilities. However, greater network connectivity and homophily may increase the prevalence of niche content and, in turn, increase polarization. In another related work, [42] analyzes content filtering in a similar two-type model within a voting framework. In their model, agents receive private signals and strategically share content based on their own partisan views and the composition of their audience on social media. Sharing decisions are strategic, whereas voting decisions are sincere and reflect Bayesian updating of beliefs about the state of the world. Different from these two models, our framework, in a multiple opinion setup, allows agents to strategically choose what to post (from a continuous opinion space) based on popularity and alignment incentives. Moreover, in a three-opinion setting, we show that the impact of strategic expression on polarization is state-dependent: polarization is amplified (dampened) when the underlying society is highly polarized (less polarized).
@@
In this framework with sticky opinions and popularity incentives, the welfare consequences are sharply state-dependent. For light topics, social media is welfare-improving because the utility gains from popularity and exposure to aligned content exceed the utility losses from exposure to misaligned yet low-stakes content. However, for intense topics, social media interaction is often harmful for some groups of agents; and for high-polarization events, intense topics create an environment in which everyone is worse off from social media interaction due to high misalignment costs. We also show that network density plays a purely welfare-amplifying role: denser networks increase exposure intensity, without altering equilibrium posting behavior. As a result, higher density magnifies welfare gains when social media interaction is welfare-improving, and amplifies welfare losses when it is welfare-reducing.
@@
Crucially, in both high- and low-polarization cases, strategic posting behavior has heterogeneous effects for different opinion holders. Relatedly, we identify a phenomenon that we call the “Popularity Trap”: for intense topics, social media activity can be utility-reducing for agents who strategically post a more popular opinion. Individually-optimal deviations by each such member cause their authentic opinion to disappear from social media, which amplifies their exposure to misaligned content, creating an equilibrium outcome where these strategic agents are worse off compared to the authentic expression benchmark. The popularity trap applies to neutral (opinionated) agents during high-(low-)polarization events.
@@
Finally, our framework is distinct from the widely studied conformity models (e.g., [7, 10, 39, 46, 17, 31, 32]). In conformity models, utility typically increases with similarity to others’ actions. These incentives generate coordination benefits: an agent benefits when others also conform to the same norm. In our framework, the direction of externalities is reversed. Agents do not desire collective convergence; rather, they prefer to enjoy higher utility from popularity, but prefer that other agents who share their authentic opinion continue to express it authentically. This generates distinct utility and welfare implications compared to conformity models, and also introduces the popularity trap.
@@ 3  Expressed Opinions and Distorted Utilities
Lemma 1 relies on a duality in equilibrium: agents’ authentic preferences are reflected in their likes on others’ posts but not necessarily in their own posts. At equilibrium, agent $i$ likes agent $j$’s post if and only if the post aligns with her own opinion, $c_{j}=b_{i}$. For any $c_{j}\neq b_{i}$, liking $j$’s post yields disutility $\omega^{d}_{i}|b_{i}-c_{j}|$, and therefore $i$ chooses not to react ($r_{ij}=0$). In contrast, when $j$’s post aligns with $i$’s opinion ($c_{j}=b_{i}$), liking provides a positive payoff $\omega^{a}_{i}$. Hence, $r_{ij}=1$ if and only if $c_{j}=b_{i}$.
@@ 4  Polarization and Welfare on Social Media
The equilibrium group sizes representing opinions b-b, 0, and +b+b on social media are denoted by $C_{-}$, $C_{0}$, and $C_{+}$, respectively. Following Lemma 1, $C_{-}+C_{0}+C_{+}=n$ holds. To avoid divisibility issues, we assume that $n$ and $G_{0}$ are even numbers. Moreover, if there exists some neutral agents who are indifferent between posting one of the opinionated views b-b and +b+b, we assume that the number of such indifferent neutral agents is always even, and the equilibrium posts of these agents are evenly split between $C_{-}$ and $C_{+}$.
@@ 7  Conclusion
Proof of Lemma 1
Proof.
Fix an equilibrium $(\mathbf{c},\mathbf{R})$ and an arbitrary agent $i\in\mathcal{N}$, and solve for the SPNE using backward induction.
Reaction stage $t=1$.
Fix $\mathbf{c}$ and consider any follower $j\in\mathcal{A}_{i}$ who observes $c_{i}$ and chooses $r_{ji}\in\{0,1\}$. Since $r_{ji}$ enters $U_{j}$ only through $R_{i}$, setting $r_{ji}=1$ increases $R_{i}$ by one unit. From (1), the induced change in $U_{j}$ is
$U_{j}(r_{ji}=1\mid\mathbf{c},\mathbf{R}_{-ji})-U_{j}(r_{ji}=0\mid\mathbf{c},\mathbf{R}_{-ji})=\omega_{j}^{a}\mathbf{1}\{c_{i}=b_{j}\}-\omega_{j}^{d}\mathbf{1}\{c_{i}\neq b_{j}\}\,|b_{j}-c_{i}|.$
where $\omega_{j}^{a}>0$ and $\omega_{j}^{d}>0$. Therefore, for $\omega_{j}^{a}>0$ and $\omega_{j}^{d}>0$, the unique best response is
$r_{ji}^{*}=1_{\{c_{i}=b_{j}\}}\qquad\text{and}\qquad r_{ji}^{*}=0_{\{c_{i}\neq b_{j}\}}.$
Therefore, given $c_{i}$, agent $i$’s equilibrium level of popularity is
$R_{i}=\sum_{j\in\mathcal{A}_{i}}r_{ji}^{*}=\sum_{j\in\mathcal{A}_{i}}1_{\{b_{j}=c_{i}\}}.$
In particular, if $c_{i}\notin\mathcal{O}$, then $r_{ji}^{*}=0_{\{c_{i}\neq b_{j}\}}$ for all $j$, hence $R_{i}=0$.
Posting stage $t=0$.
Fix the strategy profile of all agents other than $i$, and write agent $i$’s utility as
$U_{i}=H_{i}+\omega_{i}^{p}R_{i}+\omega_{i}^{a}\sum_{j\in\mathcal{N}_{i}}\mathbf{1}\{c_{j}=b_{i}\}R_{j}-\omega_{i}^{d}\sum_{j\in\mathcal{N}_{i}}\mathbf{1}\{c_{j}\neq b_{i}\}R_{j}|b_{i}-c_{j}|.$
Let $F_{i}$ collect all terms in $U_{i}$ that excludes all utility gains and losses from self-posting (i.e., terms involving only $\mathbf{c}_{-i}$ and $\mathbf{R}_{-i}$). Since $i\in\mathcal{N}_{i}$ (self-exposure), we can decompose the two exposure sums into the $j=i$ term and the $j\neq i$ terms, yielding
$U_{i}=H_{i}+F_{i}+\omega_{i}^{p}R_{i}+\omega_{i}^{a}\mathbf{1}\{c_{i}=b_{i}\}R_{i}-\omega_{i}^{d}\mathbf{1}\{c_{i}\neq b_{i}\}R_{i}\,|b_{i}-c_{i}|.$
Equivalently,
$U_{i}=H_{i}+F_{i}+R_{i}\Big(\omega_{i}^{p}+\omega_{i}^{a}\mathbf{1}\{c_{i}=b_{i}\}-\omega_{i}^{d}\mathbf{1}\{c_{i}\neq b_{i}\}|b_{i}-c_{i}|\Big).$
Now take any $c_{i}\notin\mathcal{O}$. By Step 1, $R_{i}=0$, so $U_{i}=F_{i}$. Next consider $c_{i}=b_{i}\in\mathcal{O}$. Then,
$R_{i}=\sum_{j\in\mathcal{A}_{i}}\mathbf{1}\{b_{j}=b_{i}\}=A_{i,b_{i}}>0,$
and, hence,
$U_{i,\{c_{i}=b_{i}\}}=H_{i}+F_{i}+A_{i,b_{i}}(\omega_{i}^{p}+\omega_{i}^{a})>H_{i}+F_{i}=U_{i,\{c_{i}\notin\mathcal{O}\}}.$
Thus, posting $c_{i}=b_{i}$ dominates posting any $c_{i}\notin\mathcal{O}$. Consequently, any maximizer of $U_{i}(\cdot)$ is selected from $\mathcal{O}$. Therefore, in equilibrium $c_{i}\in\mathcal{O}$ for all $i\in\mathcal{N}$.